OpenStudy (britbrat1997):

Find the y intercept of the line 5x−1y=−11

OpenStudy (theprofessor35):

Get this into slope intercept form y = mx + b

OpenStudy (welshfella):

rearrange into the form y = mx + b - the intercept will be the value of b

OpenStudy (theprofessor35):

5x−1y=−11 -5x -y = -5x - 11

OpenStudy (theprofessor35):

-y = -5x - 11 y = 5x + 11

OpenStudy (theprofessor35):

11 is in replacement of the b

OpenStudy (anonymous):

The y inetrcept is -11/5

OpenStudy (theprofessor35):

answer would be y = 11

OpenStudy (britbrat1997):

thank you guys, so would the 11 be positive or neg?

OpenStudy (britbrat1997):

it would be pos because you divided by -1 right?

OpenStudy (theprofessor35):

Yes, its positive

OpenStudy (britbrat1997):

okay thank you!

OpenStudy (anonymous):

11